# Fast Diffusion-Based Counterfactuals for Shortcut Removal and Generation

### v1

@@ 4.1 Datasets and implementation details
[…]
@@
Table 2: Complexity, time, and memory consumption. FastDiME (w/o Mask) refers to our method without mask-based inpainting. $T$ denotes the number of diffusion steps and $N$ is the number of adversarial attack update steps in ACE. Total Time (hours) is for a random subset of 1000 CelebA images, Batch Time (seconds) is calculated with a batch of 5, and GPU MU is in MiB.
Dataset Method $L_{1}$ MAD FID
CheXpert DiME 0.1107 0.7977 73.2588
FastDiME 0.0897 0.7554 61.4010
FastDiME-2 0.0946 0.7596 64.0955
FastDiME-2+ 0.0894 0.7581 62.2840
NIH DiME 0.0748 0.2536 94.3023
FastDiME 0.0546 0.2244 64.1100
FastDiME-2 0.0584 0.2386 67.8026
FastDiME-2+ 0.0536 0.2263 63.2997
ISIC DiME 0.0779 0.5240 121.5135
FastDiME 0.0631 0.3732 86.0475
FastDiME-2 0.0661 0.4428 87.4575
FastDiME-2+ 0.0621 0.3905 86.9823
4.2 Counterfactual explanations
Evaluation criteria.
Original DiME FastDiME FastDiME-2+
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(a) CheXpert: Chest X-rays with ‘pacemaker’.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(b) NIH: Chest X-rays with ‘chest drain’.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
(c) ISIC 2018: Skin lesions with ‘ruler markers’.
Original ACE $\ell_{1}$ DiME FastDiME FastDiME-2+
Figure 6: CelebA counterfactual explanations for the ‘smile’ attribute. Top: no smile normal-$\rightarrow$ smile. Bottom: smile normal-$\rightarrow$ no smile.
Results.
Tab. 1 and Tab. 3 list the results of the counterfactual explanation experiments on CelebA and medical datasets, respectively. For most of the datasets, FastDiME, and its variants outperform DiME in most of the metrics, while it remains competitive against ACE in terms of image quality. ACE achieves a significantly better COUT than our method as it is sufficient to cross the decision boundary, whereas we explicitly maximize the counterfactual class probability. This can be observed in the BKL, MAD, and FR metrics that demonstrate that our method can indeed produce samples that maximize the counterfactual class probability. Further illustrating this point, Fig. 6 shows exemplary results on CelebA, demonstrating how our methods tend to produce stronger counterfactual ‘smiling’ impressions compared to ACE. Fig. 5 shows generated shortcut counterfactuals on the medical datasets where the shortcut feature has been successfully removed from the images without significant changes to the remainder of the image. More qualitative results can be found in  Appendix B.
@@ 4.4 Shortcut learning detection pipeline
We train ResNet-18 [26] classifiers $f_{k}$ initialised with ImageNet [19] pre-trained weights on each of the datasets $\mathcal{D}_{k},k=\{100,75,50\}$ and evaluate them on the three curated test sets, $\mathrm{test}_{k}$ $\mathrm{test}_{u}$, and $\mathrm{test}^{c}_{u}$. The task labels and suspected shortcuts are shown in Tab. 4, other details including dataset sizes can be found in Appendix A.
@@
To validate the extent of shortcut learning, we evaluate each of the classifiers $f_{k}$ trained on different correlation levels $k$ with the shortcut feature $s$ using Area Under the Receiver Operating Characteristic (AUROC). To quantify the shortcut learning effect using counterfactuals, we measure the differences in model predictions resulting from counterfactually changing $s$ of the test images. To this end, we measure MAD between original images ($\mathrm{test}_{u}$) and shortcut counterfactuals ($\mathrm{test}^{c}_{u}$), and their Mean confidence Difference (MD) across two subtests $X_{s=0}$ and $X_{s=1}$ according to their true shortcut label $s$.
@@ Results.
Tab. 4 and Tab. 4 shows the results of our shortcut learning detection experiments. The significant difference between $test_{k}$ and $test_{u}$ in terms of AUROC reveals that the trained classifiers indeed learn to exploit the shortcut, with the severity of shortcut learning being correlated with the strength $k\%$ of the association in the training set. The magnitude of model prediction changes in terms of MAD and MD metrics between original test images and FastDiME shortcut counterfactuals correctly signifies stronger reliance of the models on the shortcut feature for the models trained on the more strongly biased training sets. Tab. 4, demonstrates that the proposed pipeline is indeed appropriate for detecting and quantifying shortcut learning in practice. It should be noted that the proposed shortcut detection pipeline is designed to obviate the need for diagnostic labels for new data within $test_{u}$, thereby facilitating any new databases after the trained counterfactual pipeline.
@@
$\mathrm{test}_{k}$ $\mathrm{test}_{u}$ $\mathrm{test}^{c}_{u}$ MAD MD(${}_{s=1}$) MD(${}_{s=0}$)
@@
Table 4: Results of proposed shortcut detection pipeline. The table (left) details the prediction performance on the distinct test sets, $test_{k}$ (same distribution as trainset) and $test_{u}$ (unseen balanced set), alongside the associated shortcut detection metrics. The observed performance discrepancy between $test_{k}$ and $test_{u}$ quantitatively indicates the extent to which shortcuts affect the main task’s predictive accuracy, which correlates with the shortcut detection metrics, as illustrated in the figure (right).
@@ 5 Discussion, limitations and conclusion
Original ACE $\ell_{1}$ DiME FastDiME FastDiME-2+
Figure 7: Limitations for valid counterfactuals. The first row shows a hard example (smile normal-$\rightarrow$ no smile) where all method fails to make semantically meaningful changes. In the second row (no smile normal-$\rightarrow$ smile), all DiME-based methods fail to produce a suitable sample, while ACE’s result appears natural but lacks a noticeable smile alteration. In the third row (smile normal-$\rightarrow$ no smile), we observe similar performance for ACE, while our 2-step approach offers a visually better result. The last row (smile normal-$\rightarrow$ no smile) shows an example where all methods change more than the targeted smile.
@@
As expected, our efficient gradient estimation produces denoised images $\bar{x}^{c}_{t}$ that tend to be more blurred during the very early time steps compared to the expensive fully generated images $\hat{x}^{c}_{t}$ from DiME. However, as denoised images diminish noise levels and shortcut features are often highly localized and less sensitive to blurriness, our counterfactuals do not degrade image quality. To further improve them, we apply our mask-based inpainting as long as $\bar{x}^{c}_{t}$ is not blurred, which is around $\frac{\tau}{2}$. We explored both quantitatively and quantitative the effects of our mask-based inpainting. The results of our ablation study can be found in Appendix C. Illustrative examples of the path towards the final counterfactual can be found in Appendix D and E.
@@
One important limitation of our work is that it depends on unbiased shortcut classification and image generation – image generation models have also been shown to reproduce (potentially spurious) correlations from training [42]. A counterfactual generator might thus not only change the shortcut feature but also discriminative features, e.g., the disease label. We mitigate this by training our generative shortcut classifier for medical datasets only on diseased samples, and by using inpainting to limit the spatial extent of the counterfactual changes. Our method can still result in counterfactuals that change more than the targeted shortcut (see Fig. 7). Further developing the inpainting and reducing the effect of spurious correlations on the generative shortcut model remain important avenues for future work. Note that it is often easier to train an unbiased shortcut classifier as opposed to for: Shortcut features are typically easier to predict, which is precisely what makes them vulnerable.
Using shortcut counterfactuals to mitigate shortcut learning
While we focus on generating the shortcut counterfactuals and using them to detect shortcut learning, thy also have potential to mitigate shortcut learning by augmenting the training set with samples that have the shortcut feature added or removed. Similar approaches have been successful in other domains [52, 11, 6, 46, 55], and will be the subject of future work.
@@ Conclusion.
Acknowledgements.
This work was supported by the Pioneer Centre for AI (DNRF grant nr P1), the DIREC project EXPLAIN-ME (9142-00001B), and the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606).
Supplementary Material
Appendix A Detailed implementation details
A.1 Medical datasets
Shortcut labels.
The annotations of chest drains from NIH are crow-sourced from both radiologists and non-experts [16, 50], and pacemakers are labeled using LabelMe [69]. Rulers were also annotated using LabelMe [69]. Our shorcut annotations for CheXpert-pacemaker and ISIC-Ruler marker will be publicly available.
@@ A.1 Medical datasets
Notably, the test sets for the medical cases are rather small (CheXpert-PM: 202; NIH-Drain: 164; ISIC-Ruler: 173), containing exclusively cases with existing shortcut features. This selection is more meaningful for medical applications where removing (not adding) potential shortcut features is often more useful.
A.2 FastDiME hyperparameters
For CelebA, we use the same hyperameters as in DiME [33], i.e., $\tau=60$ out of 200 re-spaced time-steps, $\lambda_{p}=30$ and $\lambda_{1}=0.05$ and $\lambda_{c}\in\{8,10,15\}$. For our mask-based inpainting, we normalize the masks, use a threshold of 0.15, and perform a dilation operation to set the mask as a square with a width and height of 5 pixels, similar to ACE [34]. For all medical datasets, we set $\tau=160$ out of 400 re-spaced time-steps, and $\lambda_{c}=1$ and $\lambda_{1}=50$ and $\lambda_{p}=0$, and use a threshold of 0.15 and a dilation of 21 for our masks. For all datasets, we set $\tau_{w}=\frac{\tau}{2}$.
A.3 Experiments on shortcut detection
@@
NIH pneumothorax chest ${}^{\dagger}$ 2771 693 1728
ISIC $*{}^{*}$ ruler markers 814 204 508
@@
  • ${}^{\dagger}$ Chest drain labels contains both non-expert annotations [16] and estimated labels by DenseNet.
  • $*{}^{*}$ Malignant lesions include melanoma, basal cell carcinoma, actinic keratosis/bowen’s disease/keratoacanthoma/squamous cell carcinoma, dermatofibromas, and vascular lesions. For more information about the diseases, please refer to [14].
Table 5: The dataset size, main task and shortcuts design for the shortcuts detection experiments.
@@
We trained the ResNet-18 [26] initialized with Image-Net [19] pre-trained weights. For medical datasets, the classifiers are trained for a maximum of 50 epochs with a learning rate of $1\times 10^{-6}$, a weight decay of 0.05, an early-stop mechanism monitoring validation loss with patience of 10 epochs, and Adam [38] optimizer. For CelebA, we trained classifiers for 20 epochs using a learning rate of $1\times 10^{-5}$, a weight decay of 0.05, and Adam [38] optimizer.
Appendix B Qualitative results
@@
Appendix C Ablation study
In this section, we quantitatively study the effect of our mask-based inpainting strategy using both CelebA and CheXpert datasets. For CelebA, we randomly select 4000 images and study the effect of our parameter $\tau_{w}$ as well as the effect of the dilation parameter. Moreover, we compare our inpainting strategy with ACE [34]’s inpating step which employs RePaint [43]. Our fully gradient-guided approach ensures the validity of counterfactuals while RePaint does not guarantee valid counterfactuals. For RePaint, we set the default hyperparameters for CelebA from ACE’s [34] original implementation, i.e., $\tau=5$ out of 50 respaced steps during inpainting. Results for the ‘smile’ and ‘age’ attributes can be found in Tab. 6 and Tab. 7, respectively. For CheXpert, we also study the effect of the threshold. Increasing the threshold results in minimal changes to the original input image. Results can be found in Tab. 8.
Method dilation $\tau_{w}$ FR FID FS MNAC CD MAD
FastDiME 30 0.990 8.19 0.76 3.09 1.98 0.82
5 45 0.988 9.76 0.74 3.27 2.08 0.82
60 0.989 11.49 0.72 3.46 1.80 0.83
30 0.997 8.68 0.73 3.43 2.01 0.85
15 45 0.997 8.61 0.73 3.41 2.10 0.85
60 0.998 8.94 0.73 3.48 1.95 0.85
FastDiME-2 5 0.996 7.48 0.77 3.01 1.94 0.83
15 0.998 8.57 0.73 3.41 2.02 0.85
FastDiME-2+ 5 30 0.990 6.99 0.78 2.88 2.08 0.82
15 30 0.998 8.64 0.73 3.43 1.93 0.85
FastDiME (w/o Mask) 0.998 10.27 0.73 3.69 1.96 0.85
FastDiME (w/o Mask) + RePaint 15 0.640 9.70 0.71 3.42 1.92 0.56
Table 6: Ablation study on a subset of CelebA. Results for the ‘smile’ attribute. FastDiME (w/o Mask) refers to our method without our mask-based inpainting. We highlight the best performances.
Method dilation $\tau_{w}$ FR FID FS MNAC CD MAD
FastDiME 30 0.987 8.83 0.74 2.65 3.84 0.69
5 45 0.989 10.73 0.72 2.83 3.79 0.70
60 0.995 12.70 0.70 3.11 3.72 0.73
30 0.998 9.78 0.71 3.11 3.78 0.75
15 45 0.999 9.80 0.70 3.12 3.91 0.75
60 0.998 9.93 0.70 3.13 3.81 0.75
FastDiME-2 5 0.995 8.08 0.75 2.62 3.77 0.71
15 0.998 9.59 0.70 3.13 3.69 0.75
FastDiME-2+ 5 30 0.988 7.49 0.76 2.43 3.99 0.69
15 30 0.999 9.40 0.71 3.10 3.62 0.75
FastDiME (w/o Mask) 0.997 11.90 0.70 3.49 3.95 0.76
FastDiME (w/o Mask) + RePaint 15 0.540 11.98 0.68 2.85 3.69 0.40
Table 7: Ablation study on a subset of CelebA. Results for the ‘age’ attribute. FastDiME (w/o Mask) refers to our method w/o our mask-based inpainting. We highlight the best performances.
Method dilation $\tau_{w}$ threshold $L_{1}$ MAD FID
@@ Appendix C Ablation study
Appendix D Denoised images for efficient gradient estimation

### v2

@@ 4.2 Counterfactual explanations
CheXpert NIH ISIC
Method $L_{1}$ MAD FID $L_{1}$ MAD FID $L_{1}$ MAD FID
DiME 0.1107 0.7977 73.2588 0.0748 0.2536 94.3023 0.0779 0.5240 121.5135
FastDiME 0.0897 0.7554 61.4010 0.0546 0.2244 64.1100 0.0631 0.3732 86.0475
FastDiME-2 0.0946 0.7596 64.0955 0.0584 0.2386 67.8026 0.0661 0.4428 87.4575
FastDiME-2+ 0.0894 0.7581 62.2840 0.0536 0.2263 63.2997 0.0621 0.3905 86.9823
(a) CheXpert: Chest X-rays with ‘pacemaker’.
(b) NIH: Chest X-rays with ‘chest drain’.
(c) ISIC 2018: Skin lesions with ‘ruler markers’.
Original Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DiME Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FastDiME Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FastDiME-2+ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Original ACE $\ell_{1}$ DiME FastDiME FastDiME-2+ Original ACE $\ell_{1}$ DiME FastDiME FastDiME-2+
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
No smile $\rightarrow$ Smile.
Smile $\rightarrow$ No smile.
Figure 6: CelebA counterfactual explanations for the ‘smile’ attribute.
Results.
Tab. 1 and Tab. 2 list the results of the counterfactual explanation experiments on CelebA and medical datasets, respectively. For most of the datasets, FastDiME, and its variants outperform DiME in most of the metrics, while it remains competitive against ACE in terms of image quality. ACE achieves a significantly better COUT than our method as it is sufficient to cross the decision boundary, whereas we explicitly maximize the counterfactual class probability. This can be observed in the BKL, MAD, and FR metrics that demonstrate that our method can indeed produce samples that maximize the counterfactual class probability. Further illustrating this point, Fig. 6 shows exemplary results on CelebA, demonstrating how our methods tend to produce stronger counterfactual ‘smiling’ impressions compared to ACE. Fig. 5 shows generated shortcut counterfactuals on the medical datasets where the shortcut feature has been successfully removed from the images without significant changes to the remainder of the image.
Table 3: Complexity, time, and memory consumption. FastDiME (w/o Mask) refers to our method without self-optimized mask scheme. $T$ denotes the number of diffusion steps and $N$ is the number of adversarial attack update steps in ACE. Total Time (hours) is for a random subset of 1000 CelebA images, Batch Time (seconds) is calculated with a batch of 5, and GPU MU is in MiB.
@@ 4.4 Shortcut learning detection pipeline
We train ResNet-18 [28] classifiers $f_{k}$ initialised with ImageNet [20] pre-trained weights on each of the datasets $\mathcal{D}_{k},k=\{100,75,50\}$ and evaluate them on the three curated test sets, $\mathrm{test}_{k}$ $\mathrm{test}_{u}$, and $\mathrm{test}^{c}_{u}$. The task labels and suspected shortcuts are shown in Tab. 4.
@@
To quantify the shortcut learning effect using counterfactuals, we measure the differences in model predictions resulting from counterfactually changing $s$ of the test images. To this end, we measure MAD between original images ($\mathrm{test}_{u}$) and shortcut counterfactuals ($\mathrm{test}^{c}_{u}$), and their Mean confidence Difference (MD) across two subtests $X_{s=0}$ and $X_{s=1}$ according to their true shortcut label $s$. Furthermore, in order to validate the extent of shortcut learning apart from the unbalanced dataset setting, we also evaluate each of the classifiers $f_{k}$ trained on different correlation levels $k$ with the shortcut feature $s$ using Area Under the Receiver Operating Characteristic (AUROC).
@@ Results.
Tab. 4 and Tab. 4 shows the results of our shortcut learning detection experiments. The magnitude of model prediction changes in terms of MAD and MD metrics between original test images and FastDiME shortcut counterfactuals correctly signifies stronger reliance of the models on the shortcut feature for the models trained on the more strongly biased training sets. The significant difference between $test_{k}$ and $test_{u}$ in terms of AUROC also indicates that the trained classifiers indeed learn to exploit the shortcut, with the severity of shortcut learning being correlated with the strength $k\%$ of the association in the training set. Tab. 4, demonstrates that the proposed pipeline is indeed appropriate for detecting and quantifying shortcut learning in practice. Notice that AUROC is only used as an indicator, as it is not a reliable measurement for shortcut learning. The difference between the AUROC of $test_{k}$ and $test_{u}$ can only tell whether the shortcut is correlated with the target label, but it can not exclude other potential reasons and cannot reveal a causal relationship. If pacemakers are highly correlated with text markers in X-rays, the AUROC results cannot determine which is the shortcut. It should also be noted that the proposed shortcut detection pipeline is designed to obviate the need for diagnostic labels for new data within $test_{u}$, thereby facilitating any new databases after the trained counterfactual pipeline.
@@
$\mathrm{test}_{k}$ $\mathrm{test}_{u}$ $\mathrm{test}^{c}_{u}$ MAD MD(s=1) MD(s=0)
@@
Table 4: Results of proposed shortcut detection pipeline. The table (left) details the prediction performance on the distinct test sets, $test_{k}$ (same distribution as trainset) and $test_{u}$ (unseen balanced set), alongside the associated shortcut detection metrics. The observed performance discrepancy between $test_{k}$ and $test_{u}$ quantitatively indicates the extent to which shortcuts affect the main task’s predictive accuracy, which correlates with the shortcut detection metrics, as illustrated in the figure (right). Although AUROC indicates the presence of shortcuts, it is not a reliable measurement, as it only shows the correlation but no causality.
@@ 5 Discussion, limitations and conclusion
Original ACE $\ell_{1}$ DiME FastDiME FastDiME-2+ Original ACE $\ell_{1}$ DiME FastDiME FastDiME-2+
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Limitations for valid counterfactuals. The left-top case shows a hard example (smile $\rightarrow$ no smile) where all method fails to make semantically meaningful changes. In the left-bottom (no smile $\rightarrow$ smile), all DiME-based methods fail to produce a suitable sample, while ACE’s result appears natural but lacks a noticeable smile alteration. In the right-top (smile $\rightarrow$ no smile), we observe similar performance for ACE, while our 2-step approach offers a visually better result. Right-bottom case (smile $\rightarrow$ no smile) shows an example where all methods change more than the targeted smile.
@@
As expected, our efficient gradient estimation produces denoised images $\bar{x}^{c}_{t}$ that tend to be more blurred during the early time steps compared to the expensive generated images $\hat{x}^{c}_{t}$ from DiME. Yet, as denoised images diminish noise levels and shortcut features are often highly localized and less sensitive to blurriness, our counterfactuals do not degrade image quality. To further improve them, we apply our self-optimized mask scheme as long as $\bar{x}^{c}_{t}$ is not blurred, which is around $\frac{\tau}{2}$. We explored quantitatively and quantitatively the effects of our self-optimized masking in the Appendix.
@@
One important limitation of our work is that it depends on unbiased shortcut classification and image generation – image generation models have also been shown to reproduce (potentially spurious) correlations from training [45]. A counterfactual generator might thus not only change the shortcut feature but also discriminative features, e.g., the disease label. We mitigate this by training our generative shortcut classifier for medical datasets only on diseased samples, and by using self-optimized masks to limit the spatial extent of the counterfactual changes. Our method can sometimes result in counterfactuals that change more than the targeted shortcut (see Fig. 7). Further development of the self-optimized masking and reducing the effect of spurious correlations on the generative shortcut model remain important avenues for future work. Note that it is often easier to train an unbiased shortcut classifier as opposed to for: Shortcut features are typically easier to predict, which is precisely what makes them vulnerable.
Using shortcut counterfactuals to mitigate shortcut learning.
While we focus on generating the shortcut counterfactuals and using them to detect shortcut learning, they also have the potential to mitigate shortcut learning by augmenting the training set with samples that have the shortcut feature added or removed. Similar approaches have been successful in other domains [55, 12, 6, 49, 58].
@@ Conclusion.
Acknowledgements
Work on this project was partially funded by the Independent Research Fund Denmark (DFF, grant number 9131-00097B), the Pioneer Centre for AI (DNRF grant nr P1), the DIREC project EXPLAIN-ME (9142-00001B), and the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (MLLS, grant NNF20OC0062606). The funding agencies had no influence on the writing of the manuscript nor on the decision to submit it for publication.
Appendix
Appendix 0.A Implementation details
0.A.1 Medical datasets
Shortcut annotations.
The annotations of chest drains from NIH are crow-sourced from both radiologists and non-experts [17, 53], and pacemakers are labeled using LabelMe [73]. Rulers are also annotated using LabelMe [73]. Our shortcut annotations for CheXpert-pacemaker and ISIC-Ruler marker are public at github.com/nina-weng/FastDiME_Med.
@@ 0.A.1 Medical datasets
Notably, the test sets for the medical cases in evaluation of counterfactual explanations are relatively small (CheXpert-PM: 202; NIH-Drain: 164; ISIC-Ruler: 173), containing exclusively cases with existing shortcut features. This selection is more meaningful for medical applications, where removing (not adding) potential shortcut features is often more useful.
0.A.2 Counterfactual loss and FastDiME hyperparameters
Counterfactual loss computation.
We leverage the denoised image to compute the counterfactual loss. For all medical datasets, $L_{1}$ loss is computed between the denoised image $\bar{x}^{c}_{t}$ and the original image $x_{0}$, while we do not include a perceptual loss. For CelebA, $L_{perc}$ is computed between the denoised image $\bar{x}^{c}_{t}$ and the original image $x_{0}$, and implemented with a pre-trained VGG-19 [65] as feature extractor, while $L_{1}$ loss is computed between the noisy image $x^{c}_{t}$ and the original image $x_{0}$, following the original implementation of DiME [35].
Hyperparameters.
For CelebA, we use the same hyperameters as in DiME [35], i.e., $\tau=60$ out of 200 re-spaced time-steps, $\lambda_{p}=30$ and $\lambda_{1}=0.05$ and $\lambda_{c}\in\{8,10,15\}$. For our self-optimized masking scheme, we normalize the masks, use a threshold of 0.15, and perform a dilation operation to set the mask as a square with a width and height of 5 pixels, similar to ACE [36]. For all medical datasets, we set $\tau=160$ out of 400 re-spaced time-steps, and $\lambda_{c}=1$ and $\lambda_{1}=50$ and $\lambda_{p}=0$, and use a threshold of 0.15 and a dilation of 21 for our masks. For all datasets, we set $\tau_{w}=\frac{\tau}{2}$.
0.A.3 Experiments on shortcut detection
@@
Table 5: Experimental design for shortcut detection. Dataset size, main task, and shortcuts design for the shortcuts detection experiments.
@@
NIH pneumothorax chest drain 2771 693 1728
ISIC malignant ruler markers 814 204 508
@@
  • Chest drain labels contains both non-expert annotations [17] and estimated labels by DenseNet.
  • Malignant lesions include melanoma, basal cell carcinoma, actinic keratosis/bowen’s disease/keratoacanthoma/squamous cell carcinoma, dermatofibromas, and vascular lesions. For more information about the diseases, please refer to [15].
@@
We train the ResNet-18 [28] classifiers initialized with Image-Net [20] pre-trained weights. For medical datasets, the classifiers are trained for a maximum of 50 epochs with a learning rate of $1\times 10^{-6}$, a weight decay of 0.05, an early-stop mechanism monitoring validation loss with patience of 10 epochs, and Adam [41] optimizer. For CelebA, we train classifiers for 20 epochs using a learning rate of $1\times 10^{-5}$, a weight decay of 0.05, and Adam [41] optimizer.
Appendix 0.B Qualitative results
@@
Method dilation parameter $\tau_{w}$ threshold $L_{1}$ MAD FID
@@
Appendix 0.C Ablation study
In this section, we quantitatively study the effect of our self-optimized masking scheme using CheXpert (see Tab. 6) CelebA (see Tab. 8 and Tab. 7). For CelebA, we randomly select 4000 images and study the effect of our parameter $\tau_{w}$ as well as the impact of the dilation parameter. Moreover, we compare our self-optimized masking scheme with ACE [36]’s inpainting step which employs RePaint [46] Our fully gradient-guided approach ensures the validity of counterfactuals while RePaint does not guarantee valid counterfactuals. For RePaint, we set the default hyperparameters for CelebA from ACE’s [36] original implementation, i.e., $\tau=5$ out of 50 respaced steps during inpainting. We also quantitatively evaluate GMD [39] including our self-optimized masking scheme for the ‘smile’ attribute in CelebA. Notably, GMD’s noisy gradients lead to both lower quality and validity compared to our proposed method. Results for the ‘smile’ and ‘age’ attributes can be found in Tab. 7 and Tab. 8, respectively. For CheXpert, we also study the effect of the threshold. Increasing the threshold results in minimal changes to the original input image. Results can be found in Tab. 6.
Table 7: Ablation study on a subset of CelebA. Results for the ‘smile’ attribute. FastDiME (w/o Mask) refers to our method without our self-optimized masking scheme. We highlight the best performances.
Method dilation parameter $\tau_{w}$ FR FID FS MNAC CD MAD
FastDiME (w/o Mask) 0.998 10.27 0.73 3.69 1.96 0.85
FastDiME (w/o Mask) + RePaint 15 0.640 9.70 0.71 3.42 1.92 0.56
GMD (/w Mask) 15 30 0.881 8.45 0.75 3.37 2.37 0.71
FastDiME 30 0.990 8.19 0.76 3.09 1.98 0.82
5 45 0.988 9.76 0.74 3.27 2.08 0.82
60 0.989 11.49 0.72 3.46 1.80 0.83
30 0.997 8.68 0.73 3.43 2.01 0.85
15 45 0.997 8.61 0.73 3.41 2.10 0.85
60 0.998 8.94 0.73 3.48 1.95 0.85
FastDiME-2 5 ✗/60 (fixed) 0.996 7.48 0.77 3.01 1.94 0.83
15 ✗/60 (fixed) 0.998 8.57 0.73 3.41 2.02 0.85
FastDiME-2+ 5 30/60 (fixed) 0.990 6.99 0.78 2.88 2.08 0.82
15 30/60 (fixed) 0.998 8.64 0.73 3.43 1.93 0.85
Table 8: Ablation study on a subset of CelebA. Results for the ‘age’ attribute. FastDiME (w/o Mask) refers to our method w/o our self-optimized masking scheme. We highlight the best performances.
Method dilation parameter $\tau_{w}$ FR FID FS MNAC CD MAD
FastDiME (w/o Mask) 0.997 11.90 0.70 3.49 3.95 0.76
FastDiME (w/o Mask) + RePaint 15 0.540 11.98 0.68 2.85 3.69 0.40
FastDiME 30 0.987 8.83 0.74 2.65 3.84 0.69
5 45 0.989 10.73 0.72 2.83 3.79 0.70
60 0.995 12.70 0.70 3.11 3.72 0.73
30 0.998 9.78 0.71 3.11 3.78 0.75
15 45 0.999 9.80 0.70 3.12 3.91 0.75
60 0.998 9.93 0.70 3.13 3.81 0.75
FastDiME-2 5 ✗/60 (fixed) 0.995 8.08 0.75 2.62 3.77 0.71
15 ✗/60 (fixed) 0.998 9.59 0.70 3.13 3.69 0.75
FastDiME-2+ 5 30/60 (fixed) 0.988 7.49 0.76 2.43 3.99 0.69
15 30/60 (fixed) 0.999 9.40 0.71 3.10 3.62 0.75
Appendix 0.D Denoised images for efficient gradient estimation